# What Breaks Knowledge Graph based RAG? Empirical Insights into Reasoning under Incomplete Knowledge

### v1

@@ Abstract
Dongzhuoran Zhou1,2, Yuqicheng Zhu2,3, Xiaxia Wang4, Hongkuan Zhou2,3, Yuan He4, Jiaoyan Chen5, Evgeny Kharlamov1,2, Steffen Staab3,7 1University of Oslo, 2Bosch Center for AI, 3University of Stuttgart, 4University of Oxford, 5The University of Manchester, 7University of Southampton dongzhuoran.zhou@de.bosch.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a widely adopted framework for enhancing large language models (LLMs) by incorporating external knowledge through a retrieve-then-generate paradigm (Khandelwal et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022; Ram et al., 2023). By conditioning the generation on the retrieved documents, RAG enables LLMs to answer questions or perform tasks using more comprehensive and up-to-date knowledge than what is stored in their parameters. To improve retrieval accuracy, enable structured reasoning and support explanation, a few recent works have proposed RAG methods based on Knowledge Graph (KG-RAG) (Han et al., 2024; Peng et al., 2024). Most of them directly use existing knowledge graphs (KGs) (Luo et al., 2024; Sun et al., 2024), while some of them construct and extend structured knowledge from unstructured documents (Fang et al., 2024). Such KG-RAG systems are expected to be well-suited for structured reasoning, where answers require synthesizing information from multiple connected facts.
Despite growing interest, current evaluation practices for KG-RAG fall short in two key ways. First, most existing benchmarks  (Yih et al., 2016; Talmor and Berant, 2018a) are constructed on top of complete KGs, where direct evidence supporting the answer is readily available. For example, given the question “Who is the brother of Justin Bieber?”, the KG contains the triple hasBrother(JustinBieber, JaxonBieber), allowing the system to answer the question without reasoning. However, real-world KGs are often incomplete, and answering such questions in practice may require reasoning over alternative paths, e.g., combining hasParent(JustinBieber, JeremyBieber) and hasChild(JeremyBieber, JaxonBieber) to infer the sibling relationship. As a result, current benchmarks do not assess whether KG-RAG methods can reason over missing knowledge or simply retrieve answers directly from explicit evidence.
Second, evaluation metrics across KG-RAG studies are often inconsistent and unreliable. Definitions of commonly used metrics such as accuracy and Hits@1 vary across studies—some equate accuracy with Hits@1, while others use accuracy as a proxy for recall, or provide vague descriptions altogether. In addition, many studies only adopt overly permissive matching criteria, treating predictions as correct if they contain the gold answer as a substring, regardless of correctness or context. This results in inflated performance estimates and makes meaningful comparison difficult.
@@ 2.1 Knowledge Graph Question Answering (KGQA)
Approaches to KGQA can be broadly categorized into the following two types: (1) Semantic parsing-based methods Yih et al. (2016); Gu et al. (2021); Ye et al. (2021) translate questions into formal executable queries (e.g., SPARQL or logical forms), which are then executed over the KG to retrieve the answer entities. These methods offer high precision and interpretability, as the reasoning process is explicitly encoded. However, they face several challenges, including difficulty in understanding semantically and syntactically complex questions, handling diverse and compositional logical forms, and managing the large search space involved in parsing multi-relation queriesLan et al. (2021). (2) Embedding-based methods Yao et al. (2019); Baek et al. (2023), by contrast, encode questions and entities into a shared vector space and rank answer candidates based on embedding similarity. While these methods are end-to-end trainable and do not require annotated logical forms, they often struggle with multi-hop reasoning Qiu et al. (2020), lack interpretability Biswas et al. (2023), and exhibit high uncertainty in their predictions Zhu et al. (2024, 2025b, 2025a).
@@ 2.2 KGQA with RAG
Recently, a new line of work has emerged that integrates large language models more tightly into the KG reasoning process. These KG-RAG methods go beyond previous KGQA approaches by coupling structured retrieval with generative reasoning capabilities of LLMs. RoG Luo et al. (2024) adopts a planning-retrieval-reasoning pipeline, where an LLM generates relation paths as plans, retrieves corresponding paths from the KG, and reasons over them for answer generation. G-Retriever He et al. (2024) retrieves subgraphs via Prize-Collecting Steiner Tree optimization and provides them to LLMs as soft prompts or text prompts for question answering. GNN-RAG Mavromatis and Karypis (2024) uses a GNN to select candidate answers and retrieves shortest paths to them, which are verbalized for LLM-based answer generation. PoG Chen et al. (2024) instead decomposes questions into subgoals and iteratively explores and self-corrects reasoning paths using a guidance-memory-reflection mechanism. ToG Sun et al. (2024) treats the LLM as an agent that interactively explores and aggregates evidence on KGs through iterative beam search, StructGPT Jiang et al. (2023) introduces an interface-based framework that enables LLMs to iteratively extract relevant evidence from structured data such as KGs and reason over the retrieved information to answer complex questions.
@@ 2.3 KGQA Benchmarks
Existing benchmarks such as WebQSP Yih et al. (2016), CWQ Talmor and Berant (2018b), and GrailQA Gu et al. (2021) are widely used to evaluate KGQA. However, during dataset construction, only question-answer pairs that yield a gold SPARQL answer on the reference KG are retained, while any unanswerable questions are discarded Yih et al. (2016); Talmor and Berant (2018b); Gu et al. (2021). This implicitly assumes that each question can be answered directly using an existing fact in the KG, overlooking the reality that KGs are often incomplete. To address this, recent works Xu et al. (2024); Zhou et al. (2025) simulate KG incompleteness by either randomly deleting triples from the whole KG or removing those along the shortest path(s) between the question and answer entities. However, these approaches have a key limitation: they cannot ensure that sufficient knowledge remains in the KG to support answering each question. This can lead to misleading evaluations, as performance drops may stem from unanswerable questions rather than limitations in the model’s reasoning ability.
@@ 2.4 LLM Reasoning Evaluation
Evaluation of LLM reasoning typically covers both logical question answering and logical consistency, with benchmarks spanning deduction, multiple-choice, proof generation, and consistency constraints such as entailment, negation, and transitivity. Datasets like LogicBench Parmar et al. (2024), ProofWriter Tafjord et al. (2020), and LogicNLI Tian et al. (2021) are widely used for these tasks, but results indicate that LLMs still struggle with robust logical reasoning Cheng et al. (2025); He et al. (2025).
@@ 3.1 Rule Mining
To ensure that questions in our benchmark require reasoning rather than direct lookup, we first identify triples that are logically inferable from other facts. We achieve this by mining high-confidence Horn rules from the original KG using the AMIE3 algorithm (Lajus et al., 2020).
AMIE3 is a widely used rule mining system designed to operate efficiently over large-scale KGs. A logical rule discovered by AMIE3 has the following form (Horn rules Horn (1951)):
@@
A substitution $\sigma$ maps every variable occurring in an atom to a entity that exists in $\mathcal{G}$. For example applying $\sigma=\{X\mapsto\texttt{Justin},Y\mapsto\texttt{Jaxon}\}$ to the atom hasSibling(X,Y) yields the grounded fact hasSibling(Justin,Jaxon). A grounding of a rule $\mathbf{B}\Rightarrow H$ is
@@
We retain only rules with high confidence and sufficient support, filtering out noisy or spurious patterns. Specifically, we run AMIE3 with a confidence threshold of 0.3, a head coverage threshold of 0.1, and a maximum rule length of 4. AMIE3 incrementally generates candidate rules via breadth-first refinement Lajus et al. (2020) and evaluates them using confidence and head coverage; only those meeting the specified thresholds are retained. Additional details on the rule generation and filtering process are provided in Appendix A.
@@ 3.2 Dataset Generation
KGs typically exhibit a "long-tail" distribution, where a small number of entities participate in a disproportionately large number of triples, while the majority appear only infrequently  (Mohamed et al., 2020; Chen et al., 2023). This imbalance can cause many generated questions to share the same answer entity, leading to biased evaluation.
@@ 3.3 Dataset Overview
To support a systematic evaluation of reasoning under knowledge incompleteness, we construct benchmark datasets based on two well established KGs: Family Sadeghian et al. (2019) and FB15k-237 Toutanova and Chen (2015). These datasets differ in size, structure, and domain coverage, enabling evaluation across both synthetic and real-world settings.
@@
  • FB15k-237 is a widely used benchmark derived from Freebase Dong et al. (2014). The graph spans 14,541 entities and 237 predicates, covering real-world domains such as people, locations, and organizations.
@@
Rule Type Family FB15k-237
Symmetry: $r(x,y)\Rightarrow r(y,x)$ 0 27
Inversion: $r_{1}(x,y)\Rightarrow r_{2}(y,x)$ 6 50
Hierarchy: $r_{1}(x,y)\Rightarrow r_{2}(x,y)$ 0 76
Composition: $r_{1}(x,y)\wedge r_{2}(y,z)\Rightarrow r_{3}(x,z)$ 56 343
Other 83 570
Total 145 1,066
@@
Each dataset instance consists of (1) a natural-language question, (2) a topic entity referenced in the question, and (3) a complete set of correct answer entities derived from the original KG. Table 2 presents representative examples from each dataset. The final question set is randomly partitioned into training, validation, and test sets using an 8:1:1 ratio. This split is applied uniformly across both datasets to ensure consistency.
@@
Table 3 summarizes the number of KG triples and generated questions in each split for both datasets, under complete and incomplete KG settings.
@@ 4.1 Evaluation Setup
Since KG-RAG models typically produce raw text sequences as output, we extract the final prediction set $\mathcal{P}_{q}$ by applying string partitioning and normalizing, following Luo et al. (2024). Details of this postprocessing step are provided in Appendix E. Without specific justification, all entities are represented by randomly assigned indices without textual labels (e.g., “Barack Obama” becomes 39) to ensure that models rely solely on knowledge from the KG rather than memorized surface forms.
@@ 5.1 Overall Performance
We evaluate six representative KG-RAG methods—RoG (Luo et al., 2024), G-Retriever (He et al., 2024), GNN-RAG (Mavromatis and Karypis, 2024), PoG (Chen et al., 2024), StructGPT (Jiang et al., 2023), and ToG (Sun et al., 2023)—on the benchmark datasets introduced in Section 3.
Figure 1: Performance comparison of KG-RAG models under incomplete (blue) and complete (red) KG settings, measured by Hits@Any (top) and F1-Score (bottom).
Figure 1 report Hits@Any and F1-scores across all methods and datasets. In most cases, both metrics drop noticeably when moving from the complete to the incomplete KG setting, highlighting the challenge posed by missing direct evidence. Precision and recall follow a similar trend (see Appendix C).
@@ 5.3 Fine-Grained Analysis by Rule Type
To enable a deeper understanding of model reasoning, we conduct a fine-grained analysis of HRR across different rule types on FB15k-237, focusing on two representative models: RoG and PoG (Figure 3). Overall, RoG exhibits greater robustness to KG incompleteness than PoG across most rule types, indicating that training-based methods can better generalize to multi-hop reasoning when direct evidence is removed. An exception arises in symmetric patterns, where PoG outperforms RoG under the incomplete setting. This is notable because symmetric relations (e.g., sibling) are inherently more robust to directionality. If a model retrieves sibling(Bieber, Jaxon), it is straightforward for an LLM to infer sibling(Jaxon, Bieber). RoG’s lower robustness in this seemingly trivial case suggests that training-based methods may overfit to relational patterns seen during training, at the cost of generalizing over simpler relations.

### v3

@@ Abstract
Dongzhuoran Zhou1,2, Yuqicheng Zhu2,3, Xiaxia Wang5, Hongkuan Zhou2,3, Yuan He4,5, Jiaoyan Chen6, Steffen Staab3,7, Evgeny Kharlamov1,2 1University of Oslo, 2Bosch Center for AI, 3University of Stuttgart, 4Amazon, 5University of Oxford, 6The University of Manchester, 7University of Southampton dongzhuoran.zhou@de.bosch.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a widely adopted framework for enhancing large language models (LLMs) by incorporating external knowledge through a retrieve-then-generate paradigm (Khandelwal2020rag; Izacard2021FiD; Borgeaud2022retro; ram2023ICRALM; zhu2025argrag). By conditioning the generation on the retrieved documents, RAG enables LLMs to answer questions or perform tasks using more comprehensive and up-to-date knowledge than what is stored in their parameters. To improve retrieval accuracy, enable structured reasoning and support explanation, recent research has pivoted toward RAG methods based on Knowledge Graph (KG-RAG) (han2024retrieval; peng2024graph). Most of them directly use existing knowledge graphs (KGs) (linhao2024rog; Jiashuo2024ToG; zhou2025gr), while some of them construct and extend structured knowledge from unstructured documents (fang2024reano). Such KG-RAG systems are expected to be well-suited for structured reasoning, where answers require synthesizing information from multiple connected facts.
Despite growing interest, current evaluation practices for KG-RAG fall short in two key ways. First, most existing benchmarks  (Yih2016webqsp; Talmor2018WSQ) are constructed on top of complete KGs, where direct evidence supporting the answer is readily available. For example, given the question “Who is the brother of Justin Bieber?”, the KG contains the triple hasBrother(JustinBieber, JaxonBieber), allowing the system to answer the question without reasoning. However, real-world KGs are often incomplete, and answering such questions in practice may require reasoning over alternative paths, e.g., combining hasParent(JustinBieber, JeremyBieber) and hasChild(JeremyBieber, JaxonBieber) to infer the sibling relationship. As a result, current benchmarks do not assess whether KG-RAG methods can reason over missing knowledge or simply retrieve answers directly from explicit evidence.
Second, evaluation protocols across KG-RAG studies lack standardization and rigor. We identify two pervasive issues: (1) ambiguous definitions, where metrics like “accuracy” fluctuate arbitrarily between Exact Match and permissive substring inclusion; and (2) implementation discrepancies, where official codebases frequently contradict their paper descriptions. For instance, several benchmarks describe using ranking-based metrics (e.g., Hits@1) but implement them as loose existence checks without any ranking. These methodological lapses result in inflated performance estimates and render cross-paper comparisons unreliable.
@@ 2.1 Knowledge Graph Question Answering (KGQA)
Approaches to KGQA can be broadly categorized into the following two types: (1) Semantic parsing-based methods Yih2016webqsp; gu2021beyond; ye2021rng translate questions into formal executable queries (e.g., SPARQL or logical forms), which are then executed over the KG to retrieve the answer entities. These methods offer high precision and interpretability, as the reasoning process is explicitly encoded. However, they face several challenges, including difficulty in understanding semantically and syntactically complex questions, handling diverse and compositional logical forms, and managing the large search space involved in parsing multi-relation querieslan2021survey. (2) Embedding-based methods yao2019kg; baek2023direct, by contrast, encode questions and entities into a shared vector space and rank answer candidates based on embedding similarity. While these methods are end-to-end trainable and do not require annotated logical forms, they often struggle with multi-hop reasoning qiu2020stepwise, lack interpretability biswas2023knowledge, and exhibit high uncertainty in their predictions zhu-etal-2024-predictive; zhu2025conformalized; zhu-etal-2025-predicate; zhu2025certainty.
@@ 2.2 KGQA with RAG
Recently, a new line of work has emerged that integrates large language models more tightly into the KG reasoning process. These KG-RAG methods go beyond previous KGQA approaches by coupling structured retrieval with generative reasoning capabilities of LLMs. RoG linhao2024rog adopts a planning-retrieval-reasoning pipeline, where an LLM generates relation paths as plans, retrieves corresponding paths from the KG, and reasons over them for answer generation. G-Retriever he2024g retrieves subgraphs via Prize-Collecting Steiner Tree optimization and provides them to LLMs as soft prompts or text prompts for question answering. GNN-RAG mavromatis2024gnn uses a GNN to select candidate answers and retrieves shortest paths to them, which are verbalized for LLM-based answer generation. PoG chen2024pog instead decomposes questions into subgoals and iteratively explores and self-corrects reasoning paths using a guidance-memory-reflection mechanism. ToG Jiashuo2024ToG treats the LLM as an agent that interactively explores and aggregates evidence on KGs through iterative beam search, StructGPT jiang2023structgpt introduces an interface-based framework that enables LLMs to iteratively extract relevant evidence from structured data such as KGs and reason over the retrieved information to answer complex questions.
@@ 2.3 KGQA Benchmarks
Existing benchmarks such as WebQSP Yih2016webqsp, CWQ talmor2018web, and GrailQA gu2021beyond are widely used to evaluate KGQA. However, during dataset construction, only question-answer pairs that yield a gold SPARQL answer on the reference KG are retained, while any unanswerable questions are discarded Yih2016webqsp; talmor2018web; gu2021beyond. This implicitly assumes that each question can be answered directly using an existing fact in the KG, overlooking the reality that KGs are often incomplete. To address this, recent works xu2024generate; zhou2025evaluating simulate KG incompleteness by either randomly deleting triples from the whole KG or removing those along the shortest path(s) between the question and answer entities. However, these approaches have a key limitation: they cannot ensure that sufficient knowledge remains in the KG to support answering each question. This can lead to misleading evaluations, as performance drops may stem from unanswerable questions rather than limitations in the model’s reasoning ability.
@@ 2.4 LLM Reasoning Evaluation
Evaluation of LLM reasoning typically covers both logical question answering and logical consistency, with benchmarks spanning deduction, multiple-choice, proof generation, and consistency constraints such as entailment, negation, and transitivity. Datasets like LogicBench parmar2024logicbench, ProofWriter tafjord2020proofwriter, and LogicNLI tian2021diagnosing are widely used for these tasks, but results indicate that LLMs still struggle with robust logical reasoning cheng2025empowering; he2025supposedly.
@@ 3.1 Rule Mining
To ensure that questions in our benchmark require reasoning rather than direct lookup, we first identify triples that are logically inferable from other facts. We achieve this by mining high-confidence Horn rules from the original KG using the AMIE3 algorithm (lajus2020fast).
AMIE3 is a widely used rule mining system designed to operate efficiently over large-scale KGs. A logical rule discovered by AMIE3 has the following form (Horn rules horn1951sentences):
@@
A substitution $\sigma$ maps every variable occurring in an atom to an entity that exists in $\mathcal{G}$. For example applying $\sigma=\{X\mapsto\texttt{Justin},Y\mapsto\texttt{Jaxon}\}$ to the atom hasSibling(X,Y) yields the grounded fact hasSibling(Justin,Jaxon). A grounding of a rule $\mathbf{B}\Rightarrow H$ is
@@
We retain only rules with high confidence and sufficient support, filtering out noisy or spurious patterns. Specifically, we run AMIE3 with a confidence threshold of 0.3, a head coverage threshold of 0.1, and a maximum rule length of 4. AMIE3 incrementally generates candidate rules via breadth-first refinement lajus2020fast and evaluates them using confidence and head coverage; only those meeting the specified thresholds are retained. Additional details on the rule generation and filtering process are provided in Appendix A.
@@ 3.2 Dataset Generation
KGs typically exhibit a "long-tail" distribution, where a small number of entities participate in a disproportionately large number of triples, while the majority appear only infrequently  (mohamed2020popularity; chen2023knowledge). This imbalance can cause many generated questions to share the same answer entity, leading to biased evaluation.
@@ 3.3 Dataset Overview
To support a systematic evaluation of reasoning under knowledge incompleteness, we construct benchmark datasets based on three well established KGs: Family sadeghian2019drum, FB15k-237 toutanova2015observed and Wikidata5m wang2021kepler. These datasets differ in size, structure, and domain coverage, enabling evaluation across both synthetic and real-world settings.
@@
  • FB15k-237 is a widely used benchmark derived from Freebase dong2014data. The graph spans 14,541 entities and 237 predicates, covering real-world domains such as people, locations, and organizations.
  • Wikidata5m is a large-scale real-world KG constructed from Wikidata (vrandevcic2014wikidata). It contains about 4.6 million entities, 822 relations, and over 20 million triples spanning diverse domains such as people, places, organizations, and events wang2021kepler.
@@
Rule Type Family FB15k-237 Wikidata5m
Symmetry: $r(x,y)\Rightarrow r(y,x)$ 0 27 19
Inversion: $r_{1}(x,y)\Rightarrow r_{2}(y,x)$ 6 50 48
Hierarchy: $r_{1}(x,y)\Rightarrow r_{2}(x,y)$ 0 76 0
Composition: $r_{1}(x,y)\wedge r_{2}(y,z)\Rightarrow r_{3}(x,z)$ 56 343 0
Other 83 570 168
Total 145 1,066 235
@@
Wikidata5m
Question: What series is 3729461 (Lumia 610) a part of?  Topic Entity: 3729461 (Lumia 610)
Answer: [1059823 (Nokia Rise)]
Direct Evidence: partOfSeries(3729461 (Lumia 610), 1059823 (Nokia Rise))
Alternative Paths: follows(2846157 (Nokia Lumia 620), 3729461 (Lumia 610)) $\wedge$ partOfSeries(2846157 (Nokia Lumia 620), 1059823 (Nokia Rise))
$\Rightarrow$ partOfSeries(3729461 (Lumia 610), 1059823 (Nokia Rise))
@@
Each dataset instance consists of (1) a natural-language question, (2) a topic entity referenced in the question, and (3) a complete set of correct answer entities derived from the original KG. Table 2 presents representative examples from each dataset. The final question set is randomly partitioned into training, validation, and test sets using an 8:1:1 ratio. This split is applied uniformly across all three datasets to ensure consistency.
@@
Table 3 summarizes the number of KG triples and generated questions in each split for all three datasets, under complete and incomplete KG settings. In Appendix H, we quantitatively verify that removing triples from the incomplete KG does not render the resulting questions unanswerable.
@@ 4 Evaluation Protocol
4.1 Standardization Gaps in Prior Work
A major obstacle in KG-RAG research is the inconsistency of evaluation protocols, which often leads to unfair and misleading comparisons. We identify two primary sources of this inconsistency:
Ambiguity in Metric Definitions.
Many studies rely on vague textual descriptions such as “whether the top-1 predicted answer is correct” for Hits@1 without specifying critical details (linhao2024rog; mavromatis2024gnn; he2024g). It is often unclear how LLM outputs are parsed, how “top-1” is defined in the absence of explicit ranking, and what constitutes correctness (e.g., exact match vs. substring inclusion). Some works omit definitions entirely, citing “previous work” without verification (Jiashuo2024ToG; chen2024pog).
Discrepancies in Implementation.
We observe significant misalignments between metric descriptions and their codebase implementations. For instance, although linhao2024rog, mavromatis2024gnn, and he2024g describe their primary metric as the ranking-based “Hits@1”, their official implementations entirely bypass ranking mechanisms. Instead, linhao2024rog and mavromatis2024gnn adopt a permissive criterion where any occurrence of the gold answer within the generated text is counted as correct. Similarly, the definition of “accuracy” fluctuates between strict Exact Match and loose substring matching across different studies (Jiashuo2024ToG; chen2024pog; jiang2023structgpt). Because many works reuse reported numbers under the assumption of comparability, these inconsistencies propagate, rendering cross-paper conclusions unreliable.
To address these issues, we adhere to the fully standardized and formally defined protocol below.
@@ 4.2 Evaluation Setup
Since KG-RAG models typically produce raw text sequences as output, we extract the final prediction set $\mathcal{P}_{q}$ by applying string partitioning and normalizing, following linhao2024rog. Details of this postprocessing step are provided in Appendix E. Without specific justification, all entities are represented by randomly assigned indices without textual labels (e.g., “Barack Obama” becomes 39) to ensure that models rely solely on knowledge from the KG rather than memorized surface forms.
@@ 5.1 Overall Performance
We evaluate six representative KG-RAG methods—RoG (linhao2024rog), G-Retriever (he2024g), GNN-RAG (mavromatis2024gnn), PoG (chen2024pog), StructGPT (jiang2023structgpt), and ToG (sun2023think)—on the benchmark datasets introduced in Section 3.
Figure 1: Performance comparison of KG-RAG models under incomplete (blue) and complete (red) KG settings, measured by Hits@Any (left) and F1-Score (right) on KGs: (a-b) Family, (c-d) FB15k-237, (e-f) Wikidata5m.
Figure 2: HHR under different KG settings. (a) Family. (b) FB15k-237. (c) Wikidata5m.
Figure 2 reports Hits@Any and F1-scores across all methods and datasets. In most cases, both metrics drop noticeably when moving from the complete to the incomplete KG setting, highlighting the challenge posed by missing direct evidence. Precision and recall follow a similar trend (see Appendix C.1).
@@ 5.3 Fine-Grained Analysis by Rule Type
To enable a deeper understanding of model reasoning, we conduct a fine-grained analysis of HHR across different rule types on FB15k-237, focusing on two representative models: RoG and PoG (Figure 3). Overall, RoG exhibits greater robustness to KG incompleteness than PoG across most rule types, indicating that training-based methods can better generalize to multi-hop reasoning when direct evidence is removed. An exception arises in symmetric patterns, where PoG outperforms RoG under the incomplete setting. This is notable because symmetric relations (e.g., sibling) are inherently more robust to directionality. If a model retrieves sibling(Bieber, Jaxon), it is straightforward for an LLM to infer sibling(Jaxon, Bieber). RoG’s lower robustness in this seemingly trivial case suggests that training-based methods may overfit to relational patterns seen during training, at the cost of generalizing over simpler relations.